Tohban Report 2016-04-06
Tohban Reports | |
---|---|
Start Date: | 30 Mar 2016 |
End Date: | 6 Apr 2016 |
Tohban: | Hazel Bain |
Tohban email: | hbain@ssl.berkeley.edu |
Next Tohban: | Juan Carlos Martinez Oliveros |
List all reports |
Solar Activity
Solar activity this week was very low.
How many GOES flares occurred?
Flares above B, C, M, X class were 14 0 0 0
RHESSI flare list contains 0 over the time range 30-Mar-16 06-Apr-16
No overlapping RHESSI flares to count, sorry
Memory Management
SSR filled up this week to ~50% due to ADP testing.
Spacecraft Status
The spacecraft is currently in the middle of an anneal procedure.
The cooldown continues to proceed slowly, but without incident. Over the past 24 hours the cold tip temperature has dropped by 0.9 degrees while the cold plate has come down 1.6 degrees.
Current conditions:
16-097-17:04:46 CP1: 162.9 K 16-097-17:04:46 CP2: 161.4 K 16-097-17:04:46 CT1: 137.6 K 16-097-17:04:46 CT2: 135.9 K 16-097-17:04:46 TST: 187.4 K
16-097-17:04:46 cp1 - ct1 = 25.43 K
16-097-17:04:46 Cryo Power: 37.0 W 16-097-17:04:46 Cryo Accel: 11.3 mG
Data Gaps
N/A
Detector issues
The detectors are currently off.
Spacecraft Management
Decimation | N/A |
HLAT Decimation | N/A |
Night time data (fronts) | N/A |
Night time data (rears) | N/A |
Require extra passes? | N/A |
Requirement for moving pointer? | N/A |
Attenuator operation | N/A |
Detector problems? | N/A |